3.1 Now suppose that a multiple-choice exam has 100 questions, each with 4 possible selections. Again, assume our beleaguered student was unable to prepare for this exam and just guessed at each question. Use MINITAB to help you answer a) to c). (8 Marks: 2 Marks for each of a), b), c) and d).)

a. What is the probability that the student gets at least one question correct?

b. What is the probability that the student gets between 15 and 35 questions (inclusive) correct?

c. What is the probability that the student passes the exam (gets 60 or more questions correct)?

d. Why does the student have a better chance of passing the exam in question 5.80 than the one in question 3.1?
Answer #1

a)( from minitab you can go to Calc tab -probability distribution -select binomial -enter 100 in trail and 0.25 in probability )

probability that the student gets at least one question correct =P(X>=0)=1-P(X<=0)=1-0.0000 =1.000000

b)

probability that the student gets between 15 and 35 questions =P(15<=X<=35)=0.990593-0.0054208=0.9851722

c) probability that the student passes the exam =P(x>=60)=1-P(X<=60)=1-1 =0.000000

d)

( please revert for question 5.80)